# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

### v1

@@ Abstract
Tabular datasets are inherently heterogeneous, posing significant challenges for developing pre-trained foundation models. The recently introduced transformer-based Tabular Prior-data Fitted Network v2 (TabPFN v2) achieves unprecedented in-context learning accuracy across multiple tabular datasets, marking a pivotal advancement in tabular foundation models. In this paper, we comprehensively evaluate TabPFN v2 on over 300 datasets, confirming its exceptional generalization capabilities on small- to medium-scale tasks. Our analysis identifies randomized feature tokens as a key factor behind TabPFN v2’s success, as they unify heterogeneous datasets into a fixed-dimensional representation, enabling more effective training and inference. To further understand TabPFN v2’s predictions, we propose a leave-one-fold-out approach, transforming TabPFN v2 into a feature extractor and revealing its capability to simplify data distributions and boost accuracy. Lastly, to address TabPFN v2’s limitations in high-dimensional, large-scale, and many-category tasks, we introduce a divide-and-conquer mechanism inspired by Chain-of-Thought prompting, enabling scalable inference. By uncovering the mechanisms behind TabPFN v2’s success and introducing strategies to expand its applicability, this study provides key insights into the future of tabular foundation models.
Machine Learning, ICML
@@ 1 Introduction
Tabular data is ubiquitous across diverse applications, including healthcare (Hyland et al., 2020), finance (Kovalerchuk & Vityaev, 2005), and scientific research (Ivanciuc et al., 2007; Hyland et al., 2020). In this format, each instance is represented as a vector of attributes, and the task of a machine learning model is to map these vector inputs to their corresponding labels (Borisov et al., 2024). Traditionally, tree-based models (Prokhorenkova et al., 2018; Chen & Guestrin, 2016) have dominated the tabular domain, but deep tabular models are increasingly showing potential to advance the field and close the performance gap (Gorishniy et al., 2021; Holzmüller et al., 2024; Ye et al., 2025).
@@
However, unlike vision and language data, where pre-trained foundation models have driven significant progress (Kirillov et al., 2023; Zhou et al., 2024), tabular data is still desperately awaiting a similar breakthrough (Somepalli et al., 2022; Hollmann et al., 2023; Onishi et al., 2023; Zhu et al., 2023; Ye et al., 2023; van Breugel & van der Schaar, 2024). A primary challenge arises from the inherent heterogeneity of tabular datasets, which often vary in dimensionality and attribute meanings, making the development of effective and versatile foundation models difficult. Additionally, there is an urgent need for such models, as many tabular datasets are small-scale—such as medical data with limited patient numbers. Training individual models from scratch for these datasets is highly sensitive to hyperparameters and struggles with insufficient data for generalization (Feurer et al., 2015; Guyon et al., 2019; Han et al., 2024).
@@
Recently, the Tabular Prior-Fitted Network v2 (TabPFN v2), proposed by Hollmann et al. (2025), has emerged as a significant step forward. Built on transformer architectures (Vaswani et al., 2017) and pre-trained on gigantic synthetic datasets (Hollmann et al., 2023, 2025), TabPFN v2 can be directly applied to downstream tasks without the need for additional tuning, effectively addressing the heterogeneity of tabular data. Specifically, TabPFN v2 takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner. When evaluated on multiple datasets covering both classification and regression tasks, TabPFN v2 achieved unprecedented accuracy compared to previous methods.
Motivated by the remarkable performance of TabPFN v2, this paper pursues four timely objectives to deepen our understanding and advance the field. First, we aim to expand the evaluation of TabPFN v2 across a diverse set of scenarios to investigate its generalizability. Second, we seek to understand the underlying mechanisms by which TabPFN v2 effectively handles data heterogeneity. Third, we strive to uncover how TabPFN v2 achieves its high prediction accuracy. Last but not least, we aim to extend TabPFN v2’s applicability beyond the comfort zone defined in Hollmann et al. (2025)—datasets with no more than 10,000 samples, 500 dimensions, and 10 classes.
TabPFN v2 consistently outperforms existing methods on small- to medium-scale datasets. We evaluate TabPFN v2 extensively on over 300 tabular datasets spanning various domains, scales, dimensionalities, and tasks (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025). Our results demonstrate TabPFN v2’s exceptional generalizability on small- to medium-scale datasets, outperforming both tree-based models and state-of-the-art deep tabular models. However, TabPFN v2’s performance diminishes on large-scale and high-dimensional datasets, indicating areas where further improvements are needed.
TabPFN v2 effectively handles data heterogeneity through randomized feature tokens. A key strength of TabPFN v2 is its ability to manage heterogeneous feature spaces, enabling applications to diverse downstream datasets without additional tuning. Given a $d$-dimensional input vector $\in\mathbb{R}^{d}$, TabPFN v2 transforms it into a $d\times k$ matrix and leverages a transformer architecture to handle variability in $d$, following (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024). The critical innovation lies in the vector-to-matrix transformation: TabPFN v2 uses a “shared” $k$-dimensional vector to lift each attribute value into a $k$-dimensional space. To differentiate attributes, a random $k$-dimensional perturbation is added to the shared vector for each attribute, consistent within a dataset but distinct across datasets [^1] This detail was identified through the supplementary material and code of (Hollmann et al., 2025).. We view this transformation as a variant of token-based tabular methods, where TabPFN v2 removes the need for dataset- and attribute-specific feature token learning, a limitation that has hindered the transferability of previous methods to diverse downstream datasets. Our analysis confirms that these feature tokens require no semantic meaning as long as they are distinct across attributes, demonstrating their efficacy.
@@
TabPFN v2 simplifies feature distributions among categories. To deepen our understanding of TabPFN v2, we investigate whether its superior in-context learning capability is accompanied by the creation of more separable feature representations. This analysis is challenging due to the distinct roles of labeled training data and unlabeled test data in TabPFN v2’s in-context learning process, resulting in non-comparable embeddings. To overcome this, we propose a leave-one-fold-out strategy to extract features from the training data that align more closely with the test data. Our findings reveal that TabPFN v2 effectively transforms tabular datasets into a nearly separable embedding space. Remarkably, training a linear model on TabPFN v2’s extracted features achieves accuracy comparable to its in-context learner, underscoring TabPFN v2’s potential as a feature encoder. This approach not only offers valuable insights into TabPFN v2 but also paves the way for broader applications and further analysis of its capabilities.
Divide-and-conquer extends TabPFN v2’s applicability. As noted in Hollmann et al. (2025) and confirmed by our study, TabPFN v2 encounters challenges with large-scale, high-dimensional, and many-category datasets. To address these limitations without extensive retraining, we introduce post-hoc divide-and-conquer strategies, reminiscent of Chain-of-Thought (CoT) prompting used in large-language models (Wei et al., 2022). For large-scale datasets, we split the training data into support and query sets, using the support set to extract TabPFN v2’s features for the query and test sets. The resulting features then form a new tabular dataset compatible with linear models, akin to representation learning strategies in the early days (Vincent et al., 2010). For high-dimensional datasets, we subsample attributes into subsets (Breiman, 2001), apply TabPFN v2 to each subset in parallel, and ensemble the predictions, similar to random forests. For a multi-class task with over 10 categories, we introduce decimal encoding inspired by Error Correcting Output Codes (ECOC) (Dietterich & Bakiri, 1995), decomposing it into multiple 10-class problems. Empirical results validate these strategies, demonstrating significant gains in TabPFN v2’s accuracy across these data regimes and highlighting the potential of advanced post-hoc methods to further expand the capabilities of tabular foundation models.
Contributions. This paper provides a timely and in-depth investigation into TabPFN v2’s strengths, limitations, and potential extensions, aiming to offer valuable insights to advance tabular foundation models. In particular:
  • We confirm TabPFN v2’s remarkable in-context learning capability across diverse small- to medium-scale tasks, positioning it as a new go-to method for these data regimes.
  • We highlight randomized feature tokens as the key driver behind TabPFN v2’s handling of heterogeneous datasets and show their applicability to other deep tabular models.
  • We introduce a simple leave-one-fold-out approach to turn TabPFN v2 into a feature encoder, enhancing its understanding and suggesting broader applications (e.g., for visualization and error diagnosis).
  • We demonstrate the effectiveness of post-hoc mechanisms to extend TabPFN v2 beyond its designated data regimes, similar to prompting strategies like CoT for LLMs.
@@ 2 Related Work
Foundation Tabular Models. Pre-trained models have revolutionized vision and language domains (Kirillov et al., 2023; Zhou et al., 2024), but their adoption in tabular data remains limited due to the inherent heterogeneity of tabular datasets. Differences in feature spaces, dimensionalities, and class distributions pose significant challenges for joint training and transferability. One solution is to utilize semantic meanings of features, as demonstrated by methods that transform tabular examples into textual representations for large language models (LLMs) (Hegselmann et al., 2023; Zhang et al., 2023a; Wang et al., 2023; Wen et al., 2024). Pre-computing feature tokens based on semantic embeddings has also been explored to improve transferability (Yan et al., 2024; Kim et al., 2024).
However, semantic meanings are often unavailable due to privacy concerns, annotation costs, or lack of describability. To address this, Ye et al. (2023) propose representing examples through similarity relationships with a fixed number of reference examples, mapping them into a latent space with a fixed dimension. The TabPFN family of models (Hollmann et al., 2023, 2025) leverages the in-context learning capabilities of transformers, directly predicting labels by positioning test instances within the context of training examples. While TabPFN v1 pads features to a fixed dimension, TabPFN v2 introduces a specialized feature tokenizer to better handle heterogeneity. Meta-learning has also been explored to obtain the weights of tabular models for downstream tasks with limited examples (Iwata & Kumagai, 2020; Bonet et al., 2024). Other pre-trained models require lightweight fine-tuning on downstream datasets to adapt to heterogeneous feature and class spaces (Liu et al., 2022; Zhang et al., 2023b; Shen et al., 2023; Zhu et al., 2023).
Variants of TabPFN. TabPFN’s success stems from its pre-training on gigantic synthetic datasets, enabling strong in-context learning performance on small-scale classification tasks (Hollmann et al., 2023). Inspired by its capabilities, various applications have been investigated, including tabular data generation (Ma et al., 2024a), anomaly detection (Ruiz-Villafranca et al., 2024), and time series forecasting (Hoo et al., 2025). The efficiency of TabPFN is highly sensitive to context size, prompting strategies to enhance scalability and performance (Feuer et al., 2023; Xu et al., 2024). Some approaches enhance TabPFN’s performance on downstream tasks by adapting the context (Thomas et al., 2024), fine-tuning specific parts of the model (Feuer et al., 2024; Liu & Ye, 2025), or even pre-training the architecture on real-world datasets (Ma et al., 2024b). Nagler (2023) analyzes TabPFN from a bias-variance perspective, shedding light on its generalization capabilities. The recently introduced TabPFN v2 extends support to regression tasks and handles larger contexts (Hollmann et al., 2025). In this paper, we conduct a comprehensive evaluation of TabPFN v2, analyze its strengths and limitations, and propose solutions to overcome scalability and applicability challenges.
3 Preliminary
We formalize the problem of learning with tabular datasets and briefly describe the core mechanism of TabPFN v2.
@@ 3 Preliminary
Figure 1: Illustration of the TabPFN v2 mechanism (Hollmann et al., 2025) for classification. $\{A_{1},\ldots,A_{d}\}$ denote $d$ attributes of the task. Training examples and test instances are combined into a tabular context set and transformed into a ${(N+1)\times d\times k}$ tensor using a combination of learnable and randomized tokens. Two types of self-attentions are applied alternately along rows (inter-sample) and columns (inter-feature). The output token corresponding to the (dummy) label of the test instance is processed through an MLP to generate a 10-class logit.
@@
Learning with a Single Tabular Dataset. A tabular dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ contains $N$ training examples, corresponding to the rows in a table. Each instance ${\bm{x}}_{i}$ is represented by $d$ features or attributes (columns in the table); its label $y_{i}$ belongs to $[C]=\{1,\ldots,C\}$ for a classification task or is a numerical value for a regression task. We assume all attributes of an instance are numerical (continuous). If categorical (discrete) attributes are present, they are transformed with ordinal/one-hot encoding beforehand. The goal of tabular machine learning is to construct a mapping $f$ from an instance to its label. Specifically, given an unseen instance ${\bm{x}}^{*}\in{\mathbb{R}}^{d}$ sampled from the same distribution as ${\mathcal{D}}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f({\bm{x}}^{*}\mid{\mathcal{D}})$. A lower prediction discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates a stronger generalization ability of $f$.
TabPFN Variants. The original TabPFN implements $f$ for classification via a Transformer-like architecture (Hollmann et al., 2023). Both training and test instances are padded to a fixed dimension $k^{\prime}$ (e.g., 100) with zeros. Then, ${\bm{x}}_{i}$ and $y_{i}$ are further mapped to $\tilde{{\bm{x}}}_{i}\in{\mathbb{R}}^{k}$ and $\tilde{{\bm{y}}}_{i}\in{\mathbb{R}}^{k}$ with linear transformations. TabPFN takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner. The task context is formulated as:
${\mathcal{C}}=\{(\tilde{{\bm{x}}}_{1}+\tilde{{\bm{y}}}_{1}),\ldots,(\tilde{{\bm{x}}}_{N}+\tilde{{\bm{y}}}_{N}),(\tilde{{\bm{x}}}^{*})\}\in{\mathbb{R}}^{k\times(N+1)}\;,$
with $N+1$ tokens of $k$ dimensions each. These tokens are processed through multiple transformer layers, which support variable-length inputs. The output token corresponding to the test instance is further mapped to a 10-class logit using a multi-layer perceptron (MLP).
The recently introduced TabPFN v2 incorporates several modifications to the original TabPFN. For example, each attribute in ${\mathcal{D}}$ is mapped to a $k$-dimensional space through linear projection. Random perturbations are added to differentiate attributes in the $k$-dimensional space, similar to a kind of positional encoding. The core idea is illustrated in Figure 1. We view this strategy as a variant of the token-based method (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024) and will discuss the construction of tokens and analyze how they handle the challenges of heterogeneous attribute spaces in section 5. Together with the mapped label embedding $\tilde{{\bm{y}}}_{i}$, an instance ${\bm{x}}_{i}$ is transformed into tokens with size $(d+1)\times k$. For a test instance ${\bm{x}}^{*}$, since its label is unknown, a dummy label (e.g., the average of training set labels) is used to generate the label embedding $\tilde{{\bm{y}}}^{*}$.
@@
Subsequently, the set of training and test instances is transformed into a ${(N+1)\times(d+1)\times k}$ tensor. Two types of self-attentions are alternately applied for inter-sample and inter-feature in-context learning. Finally, the output token corresponding to the test instance’s dummy label $\tilde{{\bm{y}}}^{*}$ is extracted and mapped to either a 10-class logit for classification or a single-value logit for regression.
@@
The weights in TabPFN v2 are pre-trained on synthetic datasets generated using structural causal models (SCMs); the checkpoint is selected based on real-world datasets. This pre-training enables TabPFN v2 to be directly applied to downstream tasks without additional tuning, effectively addressing the heterogeneity of tabular data. However, the complexity of transformers with respect to the input size limits TabPFN v2’s scalability to small and medium datasets (e.g., $N<10,000$) and its ability to handle classification with fewer than 10 classes. Please refer to Hollmann et al. (2025) for further details, including feature pre-processing, feature grouping, acceleration, and post-hoc ensembling.

### v2

# A Closer Look at TabPFN v2: Understanding Its Strengths and Extending Its Capabilities

@@ Abstract
Tabular datasets are inherently heterogeneous, presenting significant challenges for developing pre-trained foundation models. The recently introduced transformer-based Tabular Prior-data Fitted Network v2 (TabPFN v2) achieves unprecedented in-context learning performance across diverse downstream datasets, marking a pivotal advancement in tabular foundation models. In this paper, we take a closer look at TabPFN v2 to examine how it effectively handles heterogeneity and achieves high predictive accuracy, and to explore how its limitations in high-dimensional, many-category, and large-scale tasks can be mitigated. We find that TabPFN v2 can infer attribute relationships even when provided with randomized attribute token inputs, eliminating the need to explicitly learn dataset-specific attribute embeddings to address heterogeneity. We further show that TabPFN v2 can be transformed into a feature extractor, revealing its ability to construct a highly separable feature space for accurate predictions. Lastly, we demonstrate that TabPFN v2’s limitations can be addressed through a test-time divide-and-conquer strategy, enabling scalable inference without requiring re-training. By uncovering the mechanisms behind TabPFN v2’s success and introducing strategies to extend its applicability, this study offers key insights into the design of future tabular foundation models.
@@ 1 Introduction
Tabular data is ubiquitous across a wide range of applications, including healthcare [33], finance [44], and scientific research [34, 33]. In this format, each instance (e.g., a patient’s record) is represented as a vector of attributes, and the goal of a machine learning model is to map these vectors to their corresponding labels [7]. Traditionally, tree-based models [57, 12] have dominated this domain, but recent advances in deep tabular models are increasingly closing the performance gap [23, 30, 77].
@@
However, unlike vision and language domains, where pre-trained foundation models have driven significant progress [41, 82], tabular data is still desperately awaiting a similar breakthrough [64, 28, 55, 83, 78, 67]. A primary challenge arises from the inherent heterogeneity of tabular datasets, which often vary in dimensionality and attribute meanings, making the development of effective and versatile foundation models difficult. Additionally, there is an urgent need for such models, as many tabular datasets are small-scale—such as medical data with limited patient numbers. Training individual models from scratch for these datasets is highly sensitive to hyperparameter choices and often fails to generalize due to limited data [19, 25, 26].
@@
Recently, the Tabular Prior-Fitted Network v2 (TabPFN v2) [29] has emerged as a significant step forward. Built on transformer architectures [68] and pre-trained on gigantic synthetic datasets [28, 29], TabPFN v2 can be directly applied to diverse downstream tasks without additional tuning. Specifically, TabPFN v2 takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner. When evaluated across both classification and regression tasks, TabPFN v2 consistently outperforms prior tabular methods, achieving state-of-the-art accuracy.
Motivated by the remarkable performance of TabPFN v2, we aim to take a step further to understand the mechanisms behind its success [^1] We note that TabPFN v2 [29], like many recent large language models (LLMs) and foundation models, does not release its training data or training recipe. Accordingly, our focus is on understanding the properties of the released pre-trained model and exploring ways to extend its applicability.—specifically, how it effectively handles dataset heterogeneity and achieves high predictive accuracy. In addition, we investigate how to overcome its current limitations—namely, its suggested data regime of no more than 10,000 samples, 500 dimensions, and 10 classes [29]—ideally without requiring model re-training. We outline major insights as follows.
  1. 1.
    TabPFN v2 internalizes attribute token learning to handle data heterogeneity. Given an instance with $d$ attributes, TabPFN v2 transforms it into a set of fixed-dimensional tokens and uses a transformer architecture to handle variability in $d$, following [65, 23, 75]. In sharp contrast to prior methods that rely on known attribute semantics (e.g., word vectors) or learn dataset-specific attribute tokens, TabPFN v2 instead employs randomized attribute tokens—resampled at each inference. This design “syntactically” allows TabPFN v2 to be directly applied to new downstream datasets with varying dimensionalities and attribute meanings without additional tuning, but raises a fundamental question: how does it still make accurate predictions? Our analysis shows that, regardless of the randomness, TabPFN v2 can consistently infer attribute relationships through in-context learning, essentially integrating attribute token learning into the inference itself. In short, TabPFN v2 unifies representation learning and prediction within a single forward pass.
  2. 2.
@@
    TabPFN v2 can be repurposed as a feature extractor for downstream tasks. The exceptional predictive performance of TabPFN v2 suggests that it produces instance-level feature representations that are highly discriminative. However, verifying this is non-trivial, as TabPFN v2’s in-context learning mechanism assigns distinct roles to labeled training and unlabeled test instances, resulting in embeddings that are not directly comparable. To overcome this, we propose a leave-one-fold-out strategy that enables the extraction of instance features more closely aligned across training and test data. Our findings reveal that TabPFN v2 effectively maps tabular instances into a nearly linearly separable embedding space. Remarkably, training a linear model on these features yields accuracy comparable to that of TabPFN v2’s in-context learner, highlighting its potential as a powerful feature encoder. This not only offers insights into TabPFN v2’s inner workings but also opens the door to broader applications (e.g., visualization and error analysis).
  3. 3.
    Test-time divide-and-conquer effectively mitigates TabPFN v2’s limitations. As noted in [29], TabPFN v2 faces challenges when applied to high-dimensional, many-category, or large-scale datasets. Rather than resorting to model re-training, we show that these limitations can be effectively addressed through carefully designed post-hoc divide-and-conquer strategies, reminiscent of test-time scaling techniques developed for large language models [71, 52]. Empirical results show significant accuracy gains across these challenging data regimes, highlighting the potential of advanced post-hoc methods to further extend the capabilities of tabular foundation models.
Remark. This paper presents a timely and in-depth investigation into TabPFN v2, offering valuable insights for advancing tabular foundation models. While we do not propose a new architecture or training scheme, our contribution lies in the novel analysis and principled extension of TabPFN v2. This reflects a growing trend in foundation model research, where understanding, evaluating, and adapting powerful models is increasingly seen as being as impactful as designing new ones.
@@ 2 Related Work
Learning with Tabular Data.
Tabular data is prevalent across diverse fields, including healthcare, finance, and education [44, 33, 59, 2]. Tree-based models, such as XGBoost [12], LightGBM [39], and CatBoost [57], have long dominated this domain. However, recent advances in deep neural networks (DNNs) have demonstrated strong potential for tabular data [7]. Popular architectures like multi-layer perceptrons [23, 38] and Transformers [32] have been adapted to tabular tasks, alongside custom architectures designed specifically for tabular data [42, 69].
Deep tabular methods can be broadly categorized into two types. The first type directly processes raw features [30, 22, 77], sometimes incorporating feature-specific encoding strategies [21]. The second type tokenizes features, transforming an example into a set of tokens [65, 32, 60]. Comprehensive benchmarks have been developed to evaluate these methods across diverse datasets [24, 50, 76, 61], highlighting the strengths and weaknesses of deep tabular models in various scenarios.
Tabular foundation models. Pre-trained models have revolutionized the vision and language domains [41, 82], but their adoption in tabular data remains limited due to the substantial heterogeneity across datasets. Variations in attribute spaces, dimensionalities, and label distributions present significant challenges for joint training and transferability. One solution is to leverage the semantic meanings of attributes, as demonstrated by methods that convert tabular instances into textual descriptions (e.g., “Age is 25, height is 165, …”) and apply large language models for prediction [27, 80, 70, 72]. Alternatively, some approaches aim to improve transferability by pre-computing attribute tokens based on semantic embeddings [75, 40]. In practical domains such as healthcare or scientific measurement, the semantic meanings of attributes are often inaccessible due to privacy constraints, annotation costs, or a lack of describability. To address this, [78] proposed representing each instance by its similarity profile to a fixed number of nearest neighbor examples, thereby mapping it into a consistent latent space with shared dimensional semantics. The TabPFN family [28, 29] leverages the in-context learning capabilities of transformers to directly predict labels by contextualizing test instances among training examples. This strategy inspired subsequent pre-trained tabular models such as [49, 15, 58]. While TabPFN v1 pads attribute vectors to a fixed dimension, TabPFN v2 introduces a specialized attribute tokenizer to handle heterogeneous input spaces. Meta-learning has also been explored to generate model weights tailored for downstream tabular tasks with limited data [35, 6, 51]. Other pre-trained models rely on lightweight fine-tuning to adapt to variations in attribute and label spaces [45, 81, 63, 83].
Variants of TabPFN. TabPFN’s success stems from its pre-training on massive synthetic datasets, enabling strong in-context learning performance on small-scale classification tasks [28]. Motivated by its capabilities, researchers have explored a variety of applications, including tabular data generation [48], anomaly detection [62], and time series forecasting [31]. [54] provided a bias-variance analysis of TabPFN, offering insight into its generalization behavior. Another line of research focuses on improving scalability by addressing TabPFN’s sensitivity to context size [17, 74]. Further strategies to enhance downstream performance include context adaptation with nearest neighbor [66], partial fine-tuning [18, 47], pre-training on real-world datasets [49], scalable ensemble [47], and more powerful and efficient pre-training on synthetic data [58]. Most of these variants remain restricted to classification tasks due to limitations in TabPFN v1.
The recently introduced TabPFN v2 [29] extends TabPFN to support regression tasks and accommodate larger context sizes. In this paper, we conduct a comprehensive evaluation of TabPFN v2, analyze its strengths, and introduce methods to overcome its scalability and applicability challenges.
@@
Figure 1: Left: Illustration of TabPFN v2’s mechanism for binary classification [29]. $\{A_{1},\ldots,A_{d}\}$ denote $d$ attributes of the task. Training examples and a test instance are combined into a tabular context and transformed into a ${(N+1)\times(d+1)\times k}$ tensor using a combination of learnable and randomized tokens. Two types of self-attention are applied alternately across rows (inter-sample) and columns (inter-feature). The output token corresponding to the (dummy) label of the test instance is processed through an MLP to generate a 10-class logit. Right: Wilcoxon-Holm test at a significance level of 0.05 over 273 small- to medium-scale datasets. We omit the 27 datasets used to select TabPFN v2’s checkpoint from the 300 datasets in [76].
3 Background
@@
Learning with a single tabular dataset. A tabular dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ contains $N$ training examples, corresponding to the rows in a table. Each instance ${\bm{x}}_{i}$ is characterized by $d$ features or attributes (i.e., columns in the table), where $d$ typically varies across datasets. Its label $y_{i}$ belongs to $[C]=\{1,\ldots,C\}$ for a classification task or is a numerical value for a regression task. We assume that all attributes of an instance are numerical (continuous). Categorical (discrete) attributes, if present, are transformed using ordinal or one-hot encoding beforehand. The goal of tabular machine learning is to learn a mapping $f$ from instances to their labels. Specifically, given an unseen instance ${\bm{x}}^{*}\in{\mathbb{R}}^{d}$ sampled from the same distribution as ${\mathcal{D}}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f({\bm{x}}^{*}\mid{\mathcal{D}})$. A smaller discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates stronger generalizability of $f$.
TabPFN. The original TabPFN implements $f$ for classification using a transformer-like architecture [28]. Both training and test instances are first zero-padded to a fixed dimension $k^{\prime}$ (e.g., 100). Then, ${\bm{x}}_{i}$ and $y_{i}$ are linearly projected to $\tilde{{\bm{x}}}_{i}\in{\mathbb{R}}^{k}$ and $\tilde{{\bm{y}}}_{i}\in{\mathbb{R}}^{k}$, respectively. TabPFN processes both a labeled training set and an unlabeled test instance jointly, predicting the test label in an in-context learning manner. The task context is defined as ${\mathcal{C}}=\{(\tilde{{\bm{x}}}_{1}+\tilde{{\bm{y}}}_{1}),\ldots,(\tilde{{\bm{x}}}_{N}+\tilde{{\bm{y}}}_{N}),(\tilde{{\bm{x}}}^{*})\}\in{\mathbb{R}}^{(N+1)\times k}$, consisting of $N+1$ tokens, each of dimension $k$. These tokens are processed by multiple transformer layers, which accommodate variable-length inputs (i.e., variable $N$). The output token corresponding to the test instance is passed through a multi-layer perceptron (MLP) to produce a 10-class logit.
TabPFN v2. The recently proposed variant [29] introduces several key modifications. First, each of the $d$ attributes in ${\mathcal{D}}$ is embedded into a $k$-dimensional space, with random perturbations added to differentiate attributes. Together with the label embedding $\tilde{{\bm{y}}}_{i}\in\mathbb{R}^{k}$, each training instance ${\bm{x}}_{i}$ is represented by $(d+1)$ tokens with dimension $k$. For a test instance ${\bm{x}}^{*}$, where the label is unknown, a dummy label (e.g., the average label of the training set) is used to generate the label embedding $\tilde{{\bm{y}}}^{*}$.
@@
The full input to TabPFN v2—comprising the training set and the test instance—is thus represented as a tensor of shape ${(N+1)\times(d+1)\times k}$. Two types of self-attention are applied in alternation: one over samples (among the $N{+}1$ instances) and the other over attributes (among the $d{+}1$ dimensions), enabling in-context learning along both axes. Finally, the output token corresponding to the test instance’s dummy label $\tilde{{\bm{y}}}^{*}$ is extracted and mapped to a 10-class logit for classification or a single-value logit for regression. An overview of this process is illustrated in Figure 1 (left).
@@
The weights in TabPFN v2 are pre-trained on diverse synthetic datasets generated using structural causal models (SCMs), with the checkpoint selected based on real-world datasets. For additional details, including feature pre-processing, acceleration, and post-hoc ensembling, please refer to [29].
Remark. In the tabular domain, years of research into deep and foundation models have culminated in TabPFN v2 [29]—a breakthrough that, for the first time, enables deep models to consistently outperform traditional methods without fine-tuning. However, due to venue constraints, many technical details were omitted from the main paper. For example, the use of randomized tokens was documented in the supplementary material and code. In light of this, we aim to systematically analyze TabPFN v2, as we believe such a study is more impactful than proposing yet another architecture.